A car with an acceleration of 2m/s-2 starts from point P at 8:00 a.m. ten seconds later a second car with an acceleration of 3m/s-2 starts from the same point to overtake the first.


Determine the time when the second car overtakes the first.


1
Expert's answer
2021-03-21T11:37:07-0400

"s_1=\frac{a_1t^2}{2},"

"s_2=\frac{a_2(t-10)^2}{2},"

"s_1=s_2," "\implies"

"t=\frac{10}{1-\sqrt{\frac{a_1}{a_2}}}\approx54.5~s," or

8:00:54.5 a.m.